News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 10 April 2000 9.00am EDT

During last week the observing schedule was replanned to include two observations of the target Cygnus X-3 which was accepted as a Target Of Opportunity (TOO) following a major radio flare detection. Three targets were impacted: calibration observations of E0102-72.3, a portion of the observation of MCG-6-30-15 and the target RX-J1347-114. These targets are now planned for inclusion in a future schedule.

The revised schedule for last week is as follows.

---------------------------------------------

MITG J0414+0534        ACIS-S          Apr  2


B2 0755+37 ACIS-S

W3B ACIS-I

BAL 1044+3656 ACIS-S Apr 3

Q1331+170 ACIS-S

MKW 3S ACIS-I

1735-269 ACIS-S Apr 4

E0102-72.3 ACIS-I

**Cyg X-3 TOO ACIS-S/HETG

Radiation Belt

MCG-6-30-15 ACIS-S/HETG Apr 5

**Cyg X-3 TOO ACIS-S/HETG Apr 6

UM 425 ACIS-S

NGC 4406 ACIS-S

Radiation Belts Apr 7

Zeta Ori ACIS-S/HETG

4U 1538-52 ACIS-I Apr 8

RE 0751+14 HRC-S/LETG Apr 9

Zeta Ori ACIS-S/HETG

1E 0056.8-7154 ACIS-S

Radiation Belt


---------------------------------------------

A real-time procedure was run at 7.30pm EDT on 4/7 with ACIS to test a new CCD clocking mode. The new mode (dubbed the "squeegee mode") has shown marked improvement in the energy resolution of chips in the laboratory that have been radiation damaged in a similar way to the front-side illuminated chips of ACIS. The test was successful and showed results consistent with the laboratory tests. A further set of tests are planned to optimize a number of parameters before uplinking as a permanent software patch to ACIS.

The planned observations for next week are as follows.

------------------------------------------
HD 37001               ACIS-I          Apr 10

IC 443                 ACIS-I

NGC 1569               ACIS-S

MITG J2019+1127        ACIS-S          Apr 11

A2029                  ACIS-S          Apr 12

PSR B1757-24           ACIS-S

Radiation Belts

Vela X-1               ACIS-S/HETG     Apr 13

Rho Oph Core           ACIS-I

Ophiuchi Mosaic        ACIS-I          Apr 14

V1494 Aql              ACIS-I

NGC 5055               ACIS-S

A1689                  ACIS-I          Apr 15

Virgo Cluster          ACIS-I

Radiation Belts

Q0957+561              ACIS-S

NGC 4631               ACIS-S          Apr 16

------------------------------------------

The spacecraft subsystems continued to operate nominally.

Return to Status Reports